The working substance of an engine is 1.00 mol of a diatomic ideal gas. The engine operates in a cycle consisting of three steps: (1) an adiabatic expansion from an initial volume of 9.00 L to a pressure of 1.00 atm and a volume of23.4 L, (2) a compression at constant pressure to its original volume of 9.00 L, and (3) heating at constant volume to its original pressure. Find the efficiency of this cycle.
Answer #1
During the adiabatic process there is no exchange of heat , so Q1 = 0
The amount of heat entering or leaving the system during the isobaric compression is
Q2 = Cp nR?T
= (7/2)(1)(P?V)
= (7/2)(1atm)(-23.4 + 9L)
= -50.4 atmL
In adibatic process
PV^? = constant
then P1V1^? = P2V2^?
then P1 = (V2/V1)^? P2
= (23.4/9)^1.4 (1atrn)
= 3.18 atm
The amount of the heat entering or leaving the system during the constant-volume process is
Q3 = Cv nR?T
= (5/2)(1)(?PV)
= (5/2)(6 - 1)(9)
= 112.5 atmL
Therefore the workdone
W = Q1+Q2+Q3
= 0-50.4 + 112.5
= 62.1 atm L
the efficiency
e = 62.1 / 112.5
= 0.552 or 55.2%